magnitude of forceline of line charge charge

Two very long uniform lines of charge are parallel and are separated by 0.300 m. Each line of charge has charge per unit length +5.20 micro C/m.What magnitude of forcedoes one line of charge exert on a 0.0500-m section of the other line of charge?
Answer #1

0.0810N

source: back of book
Answer #2

 Electric field magnitude for an infinite wire (which is the same as a very long uniform line of charge): 

E = λ/(2π∊r) 

where 
λ is charge per unit length 
∊ is the vacuum permittivity constant of value 8.85 × 10⁻¹² m⁻³ kg⁻¹ s⁴A² 
r is the distance from wire 
L is length 

λ= Q/L 
Q = λL = 5.20*10⁻⁶ * 0.0500 = 2.6*10⁻⁷ C 

E = λ/(2π∊r) = 5.20*10⁻⁶ / (2*π*∊* 0.300) = 311715.9 N/C 

F = Q*E = 311715.9 * 2.6*10⁻⁷ = 0.0810 N

source: calculation